You are bullish on Telecom stock. The current market price is $80 per share, and you have $8,000 of your own to invest. You borrow an additional $8,000 from your broker at an interest rate of 9.0% per year and invest $16,000 in the stock.

a. What will be your rate of return if the price of Telecom stock goes up by 11% during the next year? (Ignore the expected dividend.) (Round your answer to 2 decimal places.)

Rate of return             %

b. How far does the price of Telecom stock have to fall for you to get a margin call if the maintenance margin is 30%? Assume the price fall happens immediately. (Round your answer to 2 decimal places.)

Stock price falls below            $
Answer #1

Question a

Total investment = $16,000

This implies I purchased number of shares = $16,000/$80 = 200

Final price of share = $80 * (1 + 11%) = $88.8

Gross Profit generated = 200 * (88.8 - 80) = $1,760

Now, I also need to pay 9% interest on the amount I borrowed = $8000 * 9% = $720

Hence, Net profit = $1760 - $720 = $1,040

Net rate of return = Net profit/Amount invested = $1040/$8000 = 13%

b. Let the price per share on which margin call will be made = P

So, total value of assets (or investment) = 200P

Liability or the amount loaned = $8000

Margin Rate = Equity/Assets

Margin Rate = (Assets - Liabilities)/Assets = (200P - 8000)/200P

I will receive margin call when margin rate is less than 30%

This means, margin call will be made if (200P - 8000)/200P < 0.30

(200P - 8000) < 60P

140P < 8000

P < 57.14 At a price below 57.14, I will receive a margin call.